OpenStudy (lukecrayonz):

Part 1: Find the solution to the equation 64^(3 – x) = 4^(2x). (4 points) Part 2: Using complete sentences, explain the procedure used to solve this equation. (4 points)

OpenStudy (anonymous):

\[64=4^3\] so \[64^{3-x}=4^{3(3-x)}=4^{9-3x}\] giving \[4^{9-3x}=4^{2x}\] and solve \[9-3x=2x\]

OpenStudy (anonymous):

\[64=4^3\] \[64^{3-x}=4^{3*(3-x)} = 4^{3(3-x)}\] \[4^{(9-3x)}=4^{2x}\] \[9-3x=2x\] \[9=5x\] \[x=\frac{9}5\]

OpenStudy (anonymous):

What method is this?